# Exploiting the diversity of modeling methods to probe systematic biases in strong lensing analyses

### v1

@@
Challenges inherent to high-resolution and high signal-to-noise data as well as model degeneracies can cause systematic biases in analyses of strong lens systems. In the past decade, the diversity of lens modeling methods has exploded, from being purely analytical to pixelated and non-parametric, or based on deep learning. We embrace this diversity by selecting different software packages and use them to blindly model independently simulated Hubble Space Telescope imaging data. To overcome the difficulties arising from using different codes and conventions, we use the COde-independent Organized LEns STandard (COOLEST) to store, compare and release all models in a self-consistent and human-readable manner. From an ensemble of six modeling methods, we study the recovery of the lens potential parameters and properties of the reconstructed source. In particular, we simulate and infer parameters of an elliptical power-law mass distribution with external shear for the lens while each modeling method reconstructs the source differently. We find that overall, both lens and source properties are recovered reasonably well, but systematic biases arise in all methods. Interestingly, we do not observe that a single method is significantly more accurate than others, and the amount of bias largely depends on the specific lens or source property of interest. By combining posterior distributions from individual methods using equal weights, the maximal systematic biases on lens model parameters inferred from individual models are reduced by a factor of 5.4 on average. We investigate a selection of modeling effects that partly explain the observed biases, such as the cuspy nature of the background source and the accuracy of the point spread function. This work introduces, for the first time, a generic framework to compare and ease the combination of models obtained from different codes and methods, which will be key to retain accuracy in future strong lensing analyses.
@@ 1 Introduction
Understanding the evolutionary path of galaxies over cosmic times continues to be a major challenge in astrophysics. In this context, strong gravitational lensing enables the observation of galaxies lying at different redshifts in a single observation, making it an inescapable tool to constrain galaxy evolution models. Strong gravitational lensing arises when a foreground distant galaxy—the lens, or deflector—is coincidentally aligned with a more distant background galaxy—the source—causing the appearance of multiple and magnified images of the latter. The typical redshift range for lens galaxies lies between $z_{\rm d}\sim 0.2$ and $1.5$, while source galaxies are often found between redshifts $z_{\rm s}\sim 1$ and $4$ (Oguri & Marshall 2010; Collett 2015), such that strong lensing systems can display a wide variety of galaxy morphologies and evolutionary stages.
@@
Besides galaxy evolution studies, strong lensing has several important applications in cosmology. As it is dictated by the total mass of galaxies, one can use this effect to put constraints on their dark matter (DM) halo. In particular, strong lensing data enables the separation of the baryonic and dark components of galaxies (e.g., Suyu et al. 2012; Shajib et al. 2021), and the detection of DM subhalos and other invisible masses along the line of sight (e.g., Vegetti et al. 2010; Sengül et al. 2022; Nightingale et al. 2024). When combined with time-varying sources, strong lenses can also be used to measure cosmological parameters, including the Hubble constant (Wong et al. 2020; Birrer et al. 2020; Kelly et al. 2023, e.g.,) and density parameters (e.g. with multi-plane lensing systems, Collett & Auger 2014). All these applications rely heavily on a precise characterization of both the azimuthal and radial mass profiles of lens galaxies.
@@
The central step when analyzing strong lensing data is lens modeling. The goal of this step is to model both the mass and light distribution of the lens galaxy while simultaneously reconstructing an unlensed version of the source galaxy. Lens modeling is a challenging task because inverting the lensing effect is an ill-posed problem, in particular due to known degeneracies between the lens mass distribution and the source morphology. For example, the infamous mass-sheet degeneracy (MSD, Falco et al. 1985; Schneider & Sluse 2013), a mathematically exact degeneracy between the lens mass density and the source scaling, has been studied both theoretically and practically (e.g., Birrer et al. 2016; Unruh et al. 2017; Gomer & Williams 2020; Cao et al. 2022) and can be mitigated using complementary data sets (e.g., Birrer et al. 2020; Yıldırım et al. 2023; Khadka et al. 2024). In the past twenty years, many different lens modeling techniques, ranging from analytical to pixelated techniques and neural networks, have been developed and successfully applied to real images (e.g., Warren & Dye 2003; Suyu et al. 2006; Vegetti & Koopmans 2009; Birrer et al. 2016; Nightingale et al. 2018; Galan et al. 2024). In general, these techniques have been developed with specific lensing systems, data sets and science goals in mind, and then have been extended to cover more use cases. Consequently, it is crucial to assess how these different methods compare to each other, and if their combination is warranted to improve the robustness of lensing analyses. Such a comparison enables the quantification of possible systematic biases. Additionally, if several methods leads to consistent results, those can be combined together, improving the overall precision.
@@
So far, lens modeling comparison analyses have been rare. For cluster-scale systems, a prominent work have been initiated by Meneghetti et al. (2017) by performing an extensive comparison of several modeling approaches on both simulated and real Hubble Frontier Fields clusters. However, there have not been comparable efforts for galaxy-scale strong lens systems. While some works have reanalyzed archival data with alternative modeling software packages (e.g., Birrer et al. 2016; Shajib et al. 2021), it is only recently that more quantitative comparisons between different methods have been reported (Shajib et al. 2022; Etherington et al. 2022). The Time Delay Lens Modeling Challenge (TDLMC) compared the output of different modeling and inference strategies, but focusing only on the recovery of the Hubble constant (Ding et al. 2021). Different lens modeling codes have been compared in Lefor & Futamase (2015) and Pascale et al. (2024), although using point-like multiple images rather than extended gravitational arcs as constraints. Finally, recent works from Schuldt et al. (2023b) and Gawade et al. (2024) compared neural network predictions with more classical approaches, although on with ground-based imaging data.
@@
Our goal here is to analyze imaging data similar to those obtained with the Hubble Space Telescope (HST) with different modeling and inference techniques, and study the recovery of a given set of lens parameters while reconstructing the lensed source in different ways. To our knowledge, this is the first time such a systematic and self-consistent comparison is performed. In order to maintain this novel kind of analysis tractable, we restrict the assumptions regarding the description of the lens mass distribution and properties of the data, although remaining reasonably realistic. In particular, we limit our scope to the commonly used power-law elliptical mass distribution combined with a uniform external shear. This description of the lens deflection field has proven to be a minimal but efficient prescription for modeling the observed strong lensing effect caused by large elliptical galaxies (e.g., Koopmans et al. 2006; Suyu et al. 2013; Millon et al. 2020; Shajib et al. 2021; Etherington et al. 2022; Tan et al. 2023, to cite only a few), although the simplicity of this model has known limitations (Sonnenfeld 2018; Gomer & Williams 2021; Cao et al. 2022; Etherington et al. 2023; Ruan & Keeton 2023). We also note that recent analyses of strong lenses found evidence for multipolar deviations to the elliptical power-law profile, but higher resolution than HST is warranted for robust detection (Powell et al. 2022; Stacey et al. 2024). While most lensing analyses focus on the properties of the galaxies acting as lenses, the morphology of the lensed galaxies also hold important information about galaxy formation and evolution. Current high-resolution images of strong lenses such as those from HST showcase highly structured lensed sources (e.g., Bolton et al. 2006; Garvin et al. 2022; Wang et al. 2022). Consequently, it is also crucial to assess the ability of lens modeling codes to recover the morphology of extended lensed sources.
@@
We first select different lens modeling software packages and modeling methods that are well suited to model high-resolution and high signal-to-noise (S/N) data. Since each software package typically follows different parameter definitions and model conventions, it is not possible to directly compare the modeling results. We overcome this challenge by using the COde-independent Organized LEns STandard (COOLEST, Galan et al. 2023), an open-source standard that enables storage, sharing and analysis of all lens modeling products in a uniform manner, regardless of the modeling code originally used to perform the lens modeling tasks. Included in this standard is an analysis interface allowing us to compute important quantities (e.g., effective radii and profile slopes) and visualize lens modeling results. We extensively use COOLEST in this work, both for releasing the models and data, as well as performing the analysis of the results and generating the figures.
@@ 2 Formalism of strong gravitational lensing
The main strong lensing observables are the positions and intensities of multiply lensed images of features in a background sources. These features can either be unresolved (i.e., point sources) or spatially extended. In the latter case, the lensed source appears as several arcs or as an Einstein ring surrounding the lens object and typically covering many pixels in high-resolution imaging data. We call “image plane” the (observable) plane of the sky where lensed images appear, that we place at redshift $z_{\rm d}$ of the foreground lens, also called main deflector. Observed features in the image plane are localized with a two-dimensional angular position vector $\boldsymbol{\theta}$. For conciseness, we interchangeably use the standard Cartesian coordinates $(x,y)$ to describe a position $\boldsymbol{\theta}$ in the image plane. Each feature in the lensed images has a corresponding (unobservable) angular position $\boldsymbol{\beta}$ in the “source plane” placed at the redshift of the background object $z_{\rm s}$.
The central equation in strong lensing is the lens equation, that gives the relationship between $\boldsymbol{\beta}$ and $\boldsymbol{\theta}$:
@@
where $\nabla\psi\equiv\boldsymbol{\alpha}$ is the deflection field originating from the lens potential $\psi$, the latter being a rescaled and projected version of the underlying three-dimensional gravitational potential of the lens galaxy. Usually, a more physically relevant quantity is the projected mass density of the lens, characterized by the so-called lens convergence $\kappa$ (dimensionless) obtained with a combination of second derivatives of the lens potential:
@@
As it will be useful for the discussion (Sect. 8.4), we also recall the formula of the Fermat potential, mostly relevant for time-varying sources. The Fermat potential $\Phi_{i}$ and the Fermat potential difference $\Delta\Phi_{ij}$ between a pair of lensed images $i$ and $j$ are defined as
@@
In this work we consider parametrized forms for the lens potential $\psi$ and lens convergence $\kappa$, while the surface brightness of the lensed galaxy is described following a variety of techniques. We give more details about the modeling of these different components in Sect. 3.
@@ 3 Lens modeling methods and assumptions
We consider an ensemble of “modeling methods” that each differ on two aspects: modeling assumptions and inference techniques. For instance, modeling assumptions are typically specific choices of model components (mass and light profiles, fixed or not), regularization strategies for pixelated models and necessary hyper-parameters. Inference techniques are typically minimization and sampling algorithms to obtain best-fit parameters and estimate their posterior distributions, or sequences of distinct steps (e.g., preliminary coarse and fast model fits) to converge to the best-fit solution.
In practice, a given lens modeling software package can be considered as a modeling method, as specific choices regarding the code structure, model types and optimization techniques have been made throughout its development. For this work, we selected a subset of software packages that are sufficiently different to be considered as distinct modeling methods: Lenstronomy (Birrer & Amara 2018; Birrer et al. 2021), VKL (Vernardos & Koopmans 2022), Herculens (Galan et al. 2022, 2024) and QLens (Minor et al., in prep.). Other software packages used in several published analyses so far are GLEE (Suyu & Halkola 2010; Suyu et al. 2012), PyAutoLens (Nightingale & Dye 2015; Nightingale et al. 2018, 2021), glafic (Oguri 2010) and methods from Vegetti & Koopmans (2009, and subsequent works). However, for practical reasons we only use the first set of methods, which already form a representative sample of the various modeling methods that are currently available, from fully analytical to pixelated models, with or without adaptive grids. Such methods are referred to as classical methods, in contrast to deep learning methods that we do not consider in this work (e.g., Schuldt et al. 2023a; Adam et al. 2023; Gentile et al. 2023, for some recent works), as these would require additional assumptions regarding training sets and network architectures beyond our scope. Nevertheless, we encourage future works to conduct self-consistent comparison analyses similar to ours, that involve both classical and deep learning methods (see e.g., Schuldt et al. 2023b).
@@ 3.1 Common modeling aspects
Throughout this work, we reasonably assume that the noise in the imaging data $\boldsymbol{d}$ follows a Gaussian distribution with covariance matrix $C_{d}$. In this setting, we can write the negative log-probability of the data likelihood as
@@
where $\boldsymbol{m}$ is the predicted image (i.e., the model) and $\boldsymbol{\eta}$ represents a generic vector of model parameters. We note that the data covariance matrix $C_{d}$ is assumed to be diagonal with contributions from both background noise and photon noise. In other words, we follow the widely used assumption that the noise is uncorrelated and normally distributed. With simulated data, we have access to the true data covariance matrix $C_{d}$. As in this work we do not explore the effects of inaccurate assumptions regarding noise characteristics, we give to the modelers the true matrix $C_{d}$ and use it in all lens models.
While the likelihood term in Eq. 3.1 is common to all models considered here, specific modeling assumptions such as morphological properties of the source galaxy are encoded as additional priors. Such priors priors can either be explicitly incorporated in the inference via a regularization term written as the negative of the log-prior $-\log\mathcal{P}$, or they can be implicitly defined through a choice of parametrization such as an analytical functions. Summing the likelihood and prior terms gives the full penalty or loss function $L$, which is minimized during the inference of model parameters:
@@
Modeling methods generally describe the lensing of photons from the source by casting the lens equation into a lensing operator $\boldsymbol{\mathsf{L}}$, which depends on the lens potential parameters that we denote by $\boldsymbol{\eta_{\rm\psi}}$. This operator acts on a model of the source $\boldsymbol{s}$, described by parameters $\boldsymbol{\eta_{\rm s}}$, that can be either analytical, pixelated or a representation in function basis set, as per
@@
so that we get model image $\boldsymbol{m}$ that has the same pixel size as the data, after possible downsampling by the operator $\boldsymbol{\mathsf{R}}$ and blurring by the operator $\boldsymbol{\mathsf{B}}$. The latter incorporates the effect of the point spread function (PSF) of the instrument and seeing conditions. This PSF is assumed to be known with the same spatial sampling as the data and available to all modelers. As stated in Sect. 3.6, no constraints are imposed to modelers regarding optional supersampling of ray-tracing and convolution operations. The light distribution of the lens galaxy is not modeled because we assume that the lens light has been perfectly subtracted from the data beforehand.
For modeling the lens mass distribution of the lensing galaxy—alternatively, its lens potential—we consider the commonly used power-law elliptical mass distribution (PEMD) plus external shear. This assumption is common to all the modeling techniques, namely all modelers use the same lens potential parameter vector $\boldsymbol{\eta_{\rm\psi}}$. The convergence of the PEMD is described by
@@
where $\gamma$ is the logarithmic power-law slope ($\gamma=2$ corresponding to an isothermal profile), $q_{\rm m}$ is the axis ratio and the coordinate system $(x,y)$ has been rotated by a position angle $\phi_{\rm m}$ around the lens center $(x_{0},y_{0})$. The lens potential generated by an external shear can be easily expressed in polar coordinates with the following formula:
@@
We note that the loss function of Eq. 6 has a non-linear response to lens mass parameters $\boldsymbol{\eta}_{\psi}$. However, the set of source parameters $\boldsymbol{\eta}_{\rm s}$ can formally be split into linear (light profile amplitudes) and non-linear parameters. This property is explicitly exploited by some of the modeling methods we use in this work.
@@
The modeling methods considered in this work thus mainly differ in the assumptions regarding the light distribution of the source, $\boldsymbol{s}(\boldsymbol{\eta}_{\rm s})$, which we summarize in Table 1 and describe in more details in the next subsections.

### v2

@@
Challenges inherent to high-resolution and high signal-to-noise data as well as model degeneracies can cause systematic biases in analyses of strong lens systems. In the past decade, the number of lens modeling methods has significantly increased, from purely analytical methods, to pixelated and non-parametric ones, or ones based on deep learning. We embraced this diversity by selecting different software packages and use them to blindly model independently simulated Hubble Space Telescope (HST) imaging data. To overcome the difficulties arising from using different codes and conventions, we used the COde-independent Organized LEns STandard (COOLEST) to store, compare, and release all models in a self-consistent and human-readable manner. From an ensemble of six modeling methods, we studied the recovery of the lens potential parameters and properties of the reconstructed source. In particular, we simulated and inferred parameters of an elliptical power-law mass distribution embedded in a shear field for the lens, while each modeling method reconstructs the source differently. We find that, overall, both lens and source properties are recovered reasonably well, but systematic biases arise in all methods. Interestingly, we do not observe that a single method is significantly more accurate than others, and the amount of bias largely depends on the specific lens or source property of interest. By combining posterior distributions from individual methods using equal weights, the maximal systematic biases on lens model parameters inferred from individual models are reduced by a factor of 5.4 on average. We investigated a selection of modeling effects that partly explain the observed biases, such as the cuspy nature of the background source and the accuracy of the point spread function. This work introduces, for the first time, a generic framework to compare and ease the combination of models obtained from different codes and methods, which will be key to retain accuracy in future strong lensing analyses.
@@ 1 Introduction
Understanding the evolutionary path of galaxies over cosmic times continues to be a major challenge in astrophysics. In this context, strong gravitational lensing enables the observation of galaxies lying at different redshifts in a single observation, making it an inescapable tool to constrain galaxy evolution models. Strong gravitational lensing arises when a foreground distant galaxy – the lens, or deflector – is coincidentally aligned with a more distant background galaxy – the source – causing the appearance of multiple and magnified images of the latter. The typical redshift range for lens galaxies lies between $z_{\rm d}\sim 0.2$ and $1.5$, while source galaxies are often found between redshifts $z_{\rm s}\sim 1$ and $4$ (Oguri & Marshall 2010; Collett 2015), such that strong lensing systems can display a wide variety of galaxy morphologies and evolutionary stages.
@@
Besides galaxy evolution studies, strong lensing has several important applications in cosmology. As it is dictated by the total mass of galaxies, one can use this effect to put constraints on their dark matter halo. In particular, strong lensing data enables the separation of the baryonic and dark components of galaxies (e.g., Suyu et al. 2012; Shajib et al. 2021), and the detection of dark matter subhalos and other invisible masses along the line of sight (e.g., Vegetti et al. 2010; Sengül et al. 2022; Nightingale et al. 2024). When combined with time-varying sources, strong lenses can also be used to measure cosmological parameters, including the Hubble constant ($H_{0}$ e.g., Wong et al. 2020; Birrer et al. 2020; Kelly et al. 2023) and density parameters (e.g., with multiplane lensing systems, Collett & Auger 2014). All these applications rely heavily on a precise characterization of both the azimuthal and radial mass profiles of lens galaxies.
@@
The central step when analyzing strong lensing data is lens modeling. The goal of this step is to model both the mass and light distribution of the lens galaxy, while simultaneously reconstructing an unlensed version of the source galaxy. Lens modeling is a challenging task because inverting the lensing effect is an ill-posed problem, in particular due to known degeneracies between the lens mass distribution and the source morphology. For example, the infamous mass-sheet degeneracy (MSD, Falco et al. 1985; Schneider & Sluse 2013), a mathematically exact degeneracy between the lens mass density and the source scaling, has been studied both theoretically and practically (e.g., Birrer et al. 2016; Unruh et al. 2017; Wagner 2018; Gomer & Williams 2020; Cao et al. 2022) and can be mitigated using complementary data sets (e.g., Birrer et al. 2020; Yıldırım et al. 2023; Khadka et al. 2024). In the past twenty years, many different lens modeling techniques, ranging from analytical to pixelated techniques and neural networks, have been developed and successfully applied to real images (e.g., Warren & Dye 2003; Suyu et al. 2006; Vegetti & Koopmans 2009; Birrer et al. 2016; Nightingale et al. 2018; Galan et al. 2024). In general, these techniques have been developed with specific lensing systems, data sets, and science goals in mind, and then have been extended to cover more use cases. Consequently, it is crucial to assess how these different methods compare to each other, and if their combination is warranted to improve the robustness of lensing analyses. Such a comparison enables the quantification of possible systematic biases. Additionally, if several methods lead to consistent results, those can be combined together, improving the overall accuracy.
@@
So far, lens modeling comparison analyses have been rare. For cluster-scale systems, a prominent work has been initiated by Meneghetti et al. (2017) by performing an extensive comparison of several modeling approaches on both simulated and real Hubble Frontier Fields clusters. However, there have not been comparable efforts for galaxy-scale strong lens systems. While some works have reanalyzed archival data with alternative modeling software packages (e.g., Birrer et al. 2016; Shajib et al. 2021), it is only recently that more quantitative comparisons between different methods have been reported (Shajib et al. 2022; Etherington et al. 2022). The Time Delay Lens Modeling Challenge (TDLMC) compared the output of different modeling and inference strategies, but focusing only on the recovery of $H_{0}$ (Ding et al. 2021). Different lens modeling codes have been compared in Lefor & Futamase (2015) and Pascale et al. (2024), although using point-like multiple images rather than extended gravitational arcs as constraints. Finally, recent works from Schuldt et al. (2023b) and Gawade et al. (2024) compared neural network predictions with more classical approaches, although on with ground-based imaging data.
@@
Our goal here is to analyze imaging data similar to those obtained with the Hubble Space Telescope (HST) with different modeling and inference techniques, and study the recovery of a given set of lens parameters while reconstructing the lensed source in different ways. To our knowledge, this is the first time a systematic and self-consistent comparison between a large number (six) of state-of-the-art galaxy-scale lens modeling methods has been conducted. In order to maintain this novel kind of analysis tractable, we have restricted the assumptions regarding the description of the lens mass distribution and properties of the data, although remaining reasonably realistic. In particular, we limit our scope to the commonly used power-law elliptical mass distribution embedded in a shear field. This description of the lens deflection field has proven to be a minimal but efficient prescription for modeling the observed strong lensing effect caused by large elliptical galaxies (e.g., Koopmans et al. 2006; Suyu et al. 2013; Millon et al. 2020; Shajib et al. 2021; Etherington et al. 2022; Tan et al. 2024, to cite only a few), although the simplicity of this model has known limitations (Sonnenfeld 2018; Gomer & Williams 2021; Cao et al. 2022; Etherington et al. 2024; Ruan & Keeton 2023). We also note that recent analyses of strong lenses found evidence for multipolar deviations to the elliptical power-law profile, but a higher resolution than HST is warranted for robust detection (Powell et al. 2022; Stacey et al. 2024). While most lensing analyses focus on the properties of the galaxies acting as lenses, the morphology of the lensed galaxies also hold important information about galaxy formation and evolution. Current high-resolution images of strong lenses such as those from HST showcase highly structured lensed sources (e.g., Bolton et al. 2006; Garvin et al. 2022; Wang et al. 2022). Consequently, it is also crucial to assess the ability of lens modeling codes to recover the morphology of extended lensed sources.
@@
We first selected different lens modeling software packages and modeling methods that are well suited to model high-resolution and high signal-to-noise (S/N) data. Since each software package typically follows different parameter definitions and model conventions, it is not possible to directly compare the modeling results. We overcame this challenge by using the COde-independent Organized LEns STandard (COOLEST, Galan et al. 2023), an open-source standard that enables storage, sharing, and analysis of all lens modeling products in a uniform manner, regardless of the modeling code originally used to perform the lens modeling tasks. Included in this standard is an analysis interface allowing us to compute important quantities (e.g., effective radii and profile slopes) and visualize lens modeling results. We have extensively used COOLEST in this work, both for releasing the models and data, as well as performing the analysis of the results and generating the figures.
@@ 2 Formalism of strong gravitational lensing
The main strong lensing observables are the positions and intensities of multiply lensed images of features in a background sources. These features can either be unresolved (i.e., point sources) or spatially extended. In the latter case, the lensed source appears as several arcs or as an Einstein ring surrounding the lens object and typically covering many pixels in high-resolution imaging data. We call the (observable) plane of the sky where lensed images appear the “image plane” , that we place at the redshift $z_{\rm d}$ of the foreground lens, also called the main deflector. Observed features in the image plane are localized with a two-dimensional angular position vector, $\boldsymbol{\theta}$. For conciseness, we interchangeably use the standard Cartesian coordinates $(x,y)$ to describe a position, $\boldsymbol{\theta}$, in the image plane. Each feature in the lensed images has a corresponding (unobservable) angular position, $\boldsymbol{\beta}$, in the “source plane” placed at the redshift of the background object, $z_{\rm s}$.
The central equation in gravitational lensing is the lens equation, which gives the relationship between $\boldsymbol{\beta}$ and $\boldsymbol{\theta}$:
@@
where $\nabla\psi\equiv\boldsymbol{\alpha}$ is the deflection field originating from the lens potential, $\psi$, the latter being a rescaled and projected version of the underlying three-dimensional gravitational potential of the lens galaxy. Usually, a more physically relevant quantity is the projected mass density of the lens, characterized by the so-called lens convergence, $\kappa$ (dimensionless), obtained with a combination of second derivatives of the lens potential:
@@
As it will be useful for the discussion (Sect. 8.4), we also recall the formula of the Fermat potential, mostly relevant for time-varying sources. The Fermat potential, $\Phi_{i}$, and the Fermat potential difference, $\Delta\Phi_{ij}$, between a pair of lensed images $i$ and $j$, are defined as
@@
In this work, we consider parametrized forms for the lens mass distribution, while the surface brightness of the lensed galaxy is described following a variety of techniques. We give more details about the modeling of these different components in Sect. 3.
@@ 3 Lens modeling methods and assumptions
We have considered an ensemble of “modeling methods” that each differ on two aspects: modeling assumptions and inference techniques. For instance, modeling assumptions are typically specific choices of model components (mass and light profiles, fixed or not), regularization strategies for pixelated models and necessary hyper-parameters. Inference techniques are typically minimization and sampling algorithms to obtain best-fit parameters and estimate their posterior distributions, or sequences of distinct steps (e.g., preliminary coarse and fast model fits) to converge to the best-fit solution.
In practice, a given lens modeling software package can be considered as a modeling method, as specific choices regarding the code structure, model types and optimization techniques have been made throughout its development. For this work, we selected a subset of software packages that are sufficiently different to be considered as distinct modeling methods: Lenstronomy (Birrer & Amara 2018; Birrer et al. 2021), the Very Knotty Lenser (VKL, Vernardos & Koopmans 2022), Herculens (Galan et al. 2022, 2024) and QLens (Minor et al., in prep.). Other software packages used in several published analyses so far are GLEE (Suyu & Halkola 2010; Suyu et al. 2012), PyAutoLens (Nightingale & Dye 2015; Nightingale et al. 2018, 2021), glafic (Oguri 2010) and methods from Vegetti & Koopmans (2009, and subsequent works). However, for practical reasons we only use the first set of methods, which already form a representative sample of the various modeling methods that are currently available, from fully analytical to pixelated models, with or without adaptive grids. Such methods are referred to as classical methods, in contrast to deep learning methods that we do not consider in this work (e.g., Schuldt et al. 2023a; Adam et al. 2023; Gentile et al. 2023, for some recent works), as these would require additional assumptions regarding training sets and network architectures beyond our scope. Nevertheless, we encourage future works to conduct self-consistent comparison analyses similar to ours, that involve both classical and deep learning methods (see e.g., Schuldt et al. 2023b).
@@ 3.1 Common modeling aspects
Throughout this work, we reasonably assume that the noise in the imaging data, $\boldsymbol{d}$, follows a Gaussian distribution with covariance matrix, $C_{d}$. In this setting, we can write the negative log-probability of the data likelihood as
@@
where $\boldsymbol{m}$ is the predicted image (i.e., the model) and $\boldsymbol{\eta}$ represents a generic vector of model parameters. We note that $C_{d}$ is assumed to be diagonal with contributions from both background noise and photon noise. In other words, we follow the widely used assumption that the noise is uncorrelated and normally distributed. With simulated data, we have access to the true matrix $C_{d}$. As in this work we do not explore the effects of inaccurate assumptions regarding noise characteristics, we give to the modelers the true matrix $C_{d}$ and use it in all lens models.
While the likelihood term in Eq. 3.1 is common to all models considered here, specific modeling assumptions such as morphological properties of the source galaxy are encoded as additional priors. Such priors priors can either be explicitly incorporated in the inference via a regularization term written as the negative of the log-prior $-\log\mathcal{P}$, or they can be implicitly defined through a choice of parametrization such as an analytical functions. Summing the log-likelihood and log-prior terms gives the full penalty or loss function, $L$, which is directly proportional to the log-posterior and minimized during the inference of model parameters:
@@
Modeling methods generally describe the lensing of photons from the source by casting the lens equation into a lensing operator, $\boldsymbol{\mathsf{L}}$, which depends on the lens potential parameters that we denote by $\boldsymbol{\eta_{\rm\psi}}$. This operator acts on a model of the source, $\boldsymbol{s}$, described by parameters, $\boldsymbol{\eta_{\rm s}}$, which can be either analytical, pixelated or a representation in function basis set, as per
@@
so that we get a model image, $\boldsymbol{m}$, that has the same pixel size as the data, after possible downsampling by the operator $\boldsymbol{\mathsf{R}}$ and blurring by the operator $\boldsymbol{\mathsf{B}}$. The latter incorporates the effect of the point spread function (PSF) of the instrument and seeing conditions. This PSF is assumed to be known with the same spatial sampling as the data and available to all modelers. As stated in Sect. 3.6, no constraints are imposed to modelers regarding optional supersampling of ray-tracing and convolution operations. The light distribution of the lens galaxy is not modeled because we assume that the lens light has been perfectly subtracted from the data beforehand.
For modeling the lens mass distribution of the lensing galaxy—alternatively, its lens potential—we consider the commonly used power-law elliptical mass distribution (PEMD) with an additional shear field component. This shear component has been commonly referred to as “external” shear as it captures the net effect masses external to the lens along the line of sight; we use this wording within the scope of our work but it does not necessarily holds true when modeling real systems (see e.g., Etherington et al. 2024).
The assumption of a PEMD with external shear is common to all modeling methods, namely all modelers use the same lens potential parameter vector $\boldsymbol{\eta_{\rm\psi}}$. The convergence of the PEMD is described by (Barkana 1998; Tessore & Metcalf 2015):
@@
where $\gamma$ is the logarithmic power-law slope ($\gamma=2$ corresponding to an isothermal profile), $q_{\rm m}$ is the axis ratio and the coordinate system $(x,y)$ has been rotated by a position angle $\phi_{\rm m}$ around the lens center $(x_{0},y_{0})$. The lens potential generated by an external shear can be easily expressed in polar coordinates with the following formula (e.g., Etherington et al. 2024):
@@
We note that the loss function of Eq. 6 has a non-linear response to lens mass parameters, $\boldsymbol{\eta}_{\psi}$. However, the set of source parameters, $\boldsymbol{\eta}_{\rm s}$, can formally be split into linear (light profile amplitudes) and non-linear parameters. This property is explicitly exploited by some of the modeling methods we use in this work.
@@
The modeling methods considered in this work thus mainly differ in the assumptions regarding the light distribution of the source, $\boldsymbol{s}(\boldsymbol{\eta}_{\rm s})$, which we summarize in Table 1 and describe in more detail in the next subsections.